OpenStudy (anonymous):

how to solve 2x-y=-5 4x+y=2 by using elimination

OpenStudy (anikate):

2x-y=-5 4x+y=2 then 2(2x-y=-5) 4x+y=2 then 4x-2y=-10 4x+y=2 then subtract, and 4x's get eliminated, -3y=-12 then, divide both sides by -3, y=4 there is your answer for y, now substitute it for a y in an origignal equation, 4x+4=2 then subtract both sides by 4, 4x=-2 now divide both sides by 4, x=-2/4 simplify that, x=-1/2 now your answer y=4 x=-1/2